1+tan^2a/1+cot^2a=(1-tana/1-cota)^2 solve it only from LHS

Grade:10

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Lhs

Write sqr( cot a ) = 1/ sqr tan (a)

Simplify lhs = sqr( tan a)

= Tan a * tan a

= Tan a * tan a * 1*1
= Tan * tan a * ( 1- tana)/( 1- tan a ) * ( 1-tan a)/(1+tan a)
= Take one tan a = 1/ cot a
Multiply with each bracket as per EQUATION and obtain rhs

Regards